# ShaRP: A Novel Feature Importance Framework for Ranking

### v3

@@ 5. Implementation of ShaRP
We implement the QoIs of Section 4 in the ShaRP  library. [^1] ShaRP  is available as an open-source library at https://github.com/DataResponsibly/ShaRP.  ShaRP  is implemented in Python, follows an API structure similar to scikit-learn (Buitinck et al., 2013), and is parallelized. The library can be used both to compute exact feature importance values and to approximate them to improve running times. It also implements several explanation presentation methods that are appropriate for different QoIs.
@@ 5.1. Implementation of QoIs for ranking
Algorithm 1 presents the computation of feature importance — as a vector of Shapley values $\phi(\textbf{v})$ — to QoIs that can be specified for a single item v, independently of other items. The algorithm relies on black-box access to the model that generates the outcomes (i.e., specifying an input to the model and observing the output).
Algorithm 1 takes as input a dataset $\mathcal{D}$, an item v for which the explanation is generated, the number of samples to draw $m$, and the QoI function $\iota()$ based on which to quantify feature importance.
Note that parameter $m$ is user-defined and controls the quality of approximation. Passing in the full set of items $\mathcal{D}$ and setting $m=|\mathcal{D}-1|$ results in exact computation of the Shapley values (i.e., the item v’s features are quantified against every other item in $\mathcal{D}$). Similarly, in line 3, we can restrict the maximum size of $\mathcal{S}$ and reduce the number of coalitions we use. We will provide an empirical analysis of the impact of $m$ and of maximum coalition size on performance in Section 7.2.2.
For every feature $i\in\mathcal{A}$, the algorithm computes every coalition of the remaining features $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$. For each coalition $\mathcal{S}$, it draws $m$ samples from $\mathcal{D}$. Two vectors of items are created based on this sample: $\textbf{U}_{1}$, in which feature values of the coalition vary as in U and the remaining features are fixed to their values in v; and $\textbf{U}_{2}$, in which features in the coalition and the feature of interest $i$ vary as in U, and the remaining features are fixed to their values of v. The importance of the coalition, $\phi_{i_{\mathcal{S}}}(\textbf{v})$, is computed by invoking the function $\iota()$ that computes the difference in QoI between the two vectors of items $\textbf{U}_{1}$ and $\textbf{U}_{2}$. Following (Datta et al., 2016), we define $\iota_{QoI}(\textbf{U}_{1},\textbf{U}_{2})=QoI(\textbf{U}_{1})-QoI(\textbf{U}_{2})$. This value is added to the contribution of feature $i$, $\phi_{i}(\textbf{v})$ weighted by the number of coalitions of this size (${d-1\choose|S|}$, and the total number of possible coalition sizes $d$.
Algorithm 1 Feature importance for per-item outcomes
0:  Dataset $\mathcal{D}$, item v, number of samples $m$, $\iota()$
@@
3:     for  $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$ do
@@
We use Algorithm 1 as a starting point for computing feature importance for ranked outcomes.
Score QoI
Recall that $f$ denotes the scoring function. We overload notation and denote by $f(\textbf{U})$ the computation of the score for each item in U, and return these scores as a vector. Since $\textbf{U}_{1}$ and $\textbf{U}_{2}$ are vectors of items constructed from the same vector U, they are of the same size. To quantify feature importance to an item’s score we use the $\iota()$ function and the QoI defined in the previous section, to take the mean of the (per-element) difference of $f(\textbf{U}_{1})$ and $f(\textbf{U}_{2})$.
Rank QoI
An item’s rank is computed with respect to all other items in the sample. For this reason, to quantify the impact of v’s features on its rank, we must compute the ranking of each intervened-upon instantiation of v in $\mathcal{D^{\prime}}$. This adds two steps to calculating this QoI compared to the score QoI. The item we are explaining needs to be removed from $\mathcal{D^{\prime}}$, and the score of each item $\textbf{u}_{i}\in\textbf{U}_{1}$ (and equivalently $\textbf{u}_{j}\in\textbf{U}_{2}$) needs to be compared to the scores of all items in $\mathcal{D^{\prime}}$.
@@
0:  Dataset $\mathcal{D}$, scoring function $f$, item v, $\textbf{U}_{1}$, $\textbf{U}_{2}$, number of samples $m$
@@
The computation of $\iota_{Rank}$ is summarized in Algorithm 2. We consider each a pair of corresponding (i.e., at the same position in their respective vector) samples $\textbf{u}_{1}\in\textbf{U}_{1}$ and $\textbf{u}_{2}\in\textbf{U}_{2}$, and compute $\mathcal{D}_{1}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{1}\}$ and $\mathcal{D}_{2}=\mathcal{D}\setminus\{\textbf{v}\}\cup\{\textbf{u}_{2}\}$. Next, we rank $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ on $f$, and compute the difference in ranks $\iota_{rank}=r^{-1}_{\mathcal{D}_{2},f}(\textbf{u}_{2})-r^{-1}_{\mathcal{D}_{1},f}(\textbf{u}_{1})$. Note that we are subtracting the rank of $\textbf{u}_{1}$ from the rank of $\textbf{u}_{2}$ since a lower rank corresponds to a higher (better) position in the ranking. We return the average (expected) rank difference as $\iota$.
Top-$k$ QoI
@@
Pairwise QoI
[…]
0:  items v and u, $\iota()$
1:  $\phi(\textbf{u}\succ\textbf{v})=\langle 0,\ldots,0\rangle$
2:  for $i\in\mathcal{A}$ do
3:     for $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$ do
4:        $\textbf{U}_{1}=\langle\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{u}_{\mathcal{S}}\rangle$
5:        $\textbf{U}_{2}=\langle\textbf{v}_{\mathcal{A}\setminus\{\mathcal{S}\cup i\}}\textbf{u}_{\mathcal{S}\cup i}\rangle$
6:        $\phi_{i_{\mathcal{S}}}(\textbf{u}\succ\textbf{v})=\iota(\textbf{U}_{1},\textbf{U}_{2})$
7:        $\phi_{i}(\textbf{u}\succ\textbf{v})=\phi_{i}(\textbf{u}\succ\textbf{v})+\frac{1}{d}\frac{1}{{d-1\choose|S|}}\phi_{i_{\mathcal{S}}}(\textbf{u}\succ\textbf{v})$
8:     end for
9:  end for
10:  return  $\phi(\textbf{u}\succ\textbf{v})$ that explain why u is ranked higher than v
@@ 5.2. Visualizing Feature Importance
We use three methods for visualizing experimental results.
The first, used in Figure 2, is a waterfall visualization that presents feature importance for a single item (as in (Lundberg and Lee, 2017)).
The second visualization, used in Figures 3a,  9,  10,  12, and 11, is a box-and-whiskers plot that aggregates local feature importance across ranking strata, each covering 10% of the range. The box and whiskers plots allow us to see both the median contribution of each feature and the variance in that feature’s contribution across the items in the stratum. We choose to visualize the contributions in different strata, instead of one global feature importance plot because the contributions differ per ranking stratum. Due to the ranking process, the items in each stratum have different ranges of score values. This implies that each stratum contains items with different ranges of feature values. As a result, the feature contributions per stratum can be significantly different, see Figure 9 for an example. We can further characterize the ranges of the features of each stratum if we take into account the feature distributions and the scoring function, see discussion in Section 7.3.2 for an example.
The third visualization, used in Figures 3c and 3d, is a bar graph that presents feature contributions for pairwise QoI, created from the point of view of the first item of the pair. Note that, due to the efficiency property, defined in Section 3, feature contributions add up to the difference in ranks between the two items.
6. Metrics
Recall that our focus is on explanation methods that return a numerical vector of feature attributions explaining the outcome for a given item. We can assess pair-wise explanation agreement by comparing the feature vectors of a pair of explanations. Furthermore, we can use an explanation to compute the outcome for the item being explained (e.g., its rank), and compare it to the actual observed outcome for that item. This allows us to assess fidelity of an explanation. Below, we describe explanation agreement and fidelity metrics and also explain how these primitives can be aggregated to assess sensitivity and fidelity of an explanation method, and to quantify inter-method explanation agreement.
Notation
In Section 3, we have been using $\phi(\textbf{v})$ to represent the vector of feature weights, computed using Shapley values. We generalize our definition here to $g(\textbf{v})$ to represent the output of any feature-based explanation method $g$, regardless of whether it consists of Shapley values or of some other quantification of feature importance, with weights. For all methods we consider, $g(\textbf{v})$ is a vector of numerical contributions of each feature towards the outcome for item v.
6.1. Fidelity Metrics
Explanation Fidelity
A useful property of feature-based explanations is that the actual outcome can be computed from them. For Shapley-value-based explanations, this follows from the efficiency property of Shapley values, see Section 3. Fidelity measures how well the explanation $g(\textbf{v})$ matches the model prediction $f(\textbf{v})$ being explained, see (Chowdhury et al., 2023; Molnar, 2020). SHAP and LIME explanations can be used to compute an item’s score (score QoI in our terminology), with feature importance indicating the displacement due to that feature from the mean score, either positively or negatively. ShaRP  explanations can be used to compute the outcome for all supported QoIs, including score, rank, top-$k$, and pairwise.
For QoIs that concern a single item, namely, score, rank, and top-$k$, we compute fidelity of explanation $g$ for item v as:
Here, $\mathsf{QoI}(\textbf{v})$ returns the value of the quantity of interest (i.e., the outcome being explained by $g$), such as v’s score, rank, or presence in the top-$k$, while $g(i,\textbf{v})$ is the contribution of the i-th feature of $v$. Finally, $Z$ is the normalizer set to the maximum distance between a pair of outcomes for the given dataset $\mathcal{D}$ and ranker $f$ (omitted to simplify notation), and for the specified QoI. Note that, for pairwise QoI, fidelity $F(g,\textbf{u}\succ\textbf{v})=1$ if $u$ is ranked higher than $u$ and if $g$ predicts that relative order among the items, and is 00 otherwise.
Method Fidelity
For pairwise QoI, we compute $F(g,\mathcal{D})$ as the expectation of $F(g,\textbf{u}\succ\textbf{v})$ over all pairs of distinct items $\textbf{u},\textbf{v}\in\mathcal{D}$.
6.2. Agreement Metrics
Explanation Agreement
[…]
Method Agreement
Method Sensitivity
@@ 7. Experimental Evaluation of ShaRP
7.1. Datasets

### v4

@@ 7. The  ShaRP  library
ShaRP  is implemented in Python, follows an API structure similar to scikit-learn (Buitinck et al., 2013), and is parallelized. The library can be used both to compute exact feature importance values and to approximate them to improve running times.
@@ 7.1. Implementation of QoIs for ranking
We provide Algorithm 1 to showcase the flexibility of ShaRP. Using this implementation, we can 1) easily switch between QoIs, 2) calculate both the marginal and baseline Shapley values, and 3) approximate the Shapley values for efficiency. The algorithm relies on black-box access to the model that generates the outcome (i.e., specifying an input to the model and observing the outcome used in the QoI). Specifically, algorithm 1 takes as input a dataset $\mathcal{D}$, a reference set $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ from which we draw the samples, an item v for which the explanation is generated, the number of samples to draw $m$, the maximum coalition size $c$, and the iota function $\iota()$ (Equation 5) based on which to quantify feature importance. To change the QoI we change the input $\iota()$ function. To switch to the pairwise baseline method, we set $\mathcal{D}^{\prime}=\textbf{u}$ and $m=1$, where u is the baseline sample we want to compare v to. And finally, to approximate the feature importance, we can use the parameters $m$ and $c$. Passing in the full set of items as the reference set $\mathcal{D}^{\prime}=\mathcal{D}$, and setting $m=|\mathcal{D}-1|$ and $c=|\mathcal{A}-1|$ results in the exact computation of the Shapley values (i.e., the item v’s features are quantified against every other item in $\mathcal{D}$ using all possible coalitions for all features except $i$). Because we are computing the rank of each item against the entire dataset $\mathcal{D}$, we need to provide the dataset in the input along with the reference set. We will provide an empirical analysis of the impact of $m$ and of maximum coalition size on performance in Section 8.3.2.
We now describe the algorithm for the marginal exact computation which encompasses all cases as we explained above. By definition, when computing feature importance, Shapley values use all possible coalitions of features and all items in the dataset. We refer to this as the exact computation of local feature-based explanations. We choose to explicitly provide the construction of the random variable U in lines 4-7 of Algorithm 1 for illustrative purposes. For every feature $i\in\mathcal{A}$, the algorithm computes every coalition of the remaining features $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$. For each coalition $\mathcal{S}$, it draws $m=|\mathcal{D}-1|$ samples from $\mathcal{D}$. Two random vectors of items are created based on these samples: $\textbf{U}_{1}$, in which feature values of the coalition vary as in U and the remaining features are fixed to their values in v; and $\textbf{U}_{2}$, in which features in the coalition and the feature of interest $i$ vary as in U, and the remaining features are fixed to their values of v. The importance of the coalition, $\phi_{i_{\mathcal{S}}}(\textbf{v})$, is computed by invoking the function $\iota()$ that computes the difference in QoI between the two vectors of items $\textbf{U}_{1}$ and $\textbf{U}_{2}$. This value is added to the contribution of feature $i$, $\phi_{i}(\textbf{v})$ weighted by the number of coalitions of this size (${d-1\choose|S|}$, and the total number of possible coalition sizes $d$.
Algorithm 1 Local feature importance using ShaRP
0:  Dataset $\mathcal{D}^{\prime}$, reference set $\mathcal{D}^{\prime}$, item v, number of samples $m$, maximum coalition size c, $\iota()$
@@
3:     for  $\mathcal{S}\subseteq\mathcal{A}\setminus\{i\}$ and $|\mathcal{S}|\leq c$ do
@@
Evaluating the $\iota()$ function, is straightforward for the score QoI but not for the ranking-specific QoIs. Specifically, for the score QoI, using the definition in Section 4, we take the mean of the (per-element) difference of $f(\textbf{U}_{1})$ and $f(\textbf{U}_{2})$. However, this is not the case for ranking-specific QoIs. The rank of an item is computed with respect to all other items in the sample. This adds two steps to calculating the rank QoI compared to the score QoI. The item we are explaining needs to be removed from $\mathcal{D^{\prime}}$, and the score of each item $\textbf{u}_{1}\in\textbf{U}_{1}$ (and equivalently $\textbf{u}_{2}\in\textbf{U}_{2}$) needs to be compared to the scores of all items in $\mathcal{D^{\prime}}$. The computation of $\iota_{Rank}$ is summarized in Algorithm 2.
@@
0:  Dataset $\mathcal{D}^{\prime}$, scoring function $f$, item v, $\textbf{U}_{1}$, $\textbf{U}_{2}$, number of samples $m$
@@ 7.2. Visualizing Feature Importance
We use three methods for visualizing experimental results. The first, used in Figure 2, is a waterfall visualization that presents feature importance for a single item (as in (Lundberg and Lee, 2017)). The second visualization, used in Figures 34a5, and  10, is a box-and-whiskers plot that aggregates local feature importance across ranking strata, each covering 10% of the range. This visualization allows us to see both the median contribution of each feature and the variance in that feature’s contribution across the items in the stratum. The third visualization, used in Figures 4c and 4d, is a bar graph that presents feature contributions for the pairwise method. The visualization is created from the point of view of the first item of the pair.
@@ 8. Experimental Evaluation of ShaRP
We evaluate the performance of ShaRP  and compare it to other local feature importance methods, using several real and synthetic datasets, with the corresponding ranking tasks. Dataset properties, along with ranker type (score-based or learned) are summarized in Table 1 and described below. We present the dataset information and plot the relationship between score and rank for score-based ranking tasks in Appendix A.
Table 1. Datasets used in our experimental study. S stands for score-based ranked task and LtR for learn-to-rank.
name source # tuples # features task
Tennis (ATP) (Joao Fonseca, 2024b) 86 6 S
CS Rankings (CSR) (Berger, 2023) 189 5 S
Times Higher Education (THE) (Joao Fonseca, 2024a) 1,397 5 S
Synthetic (SYN) here 2,000 2 or 3 S
Moving company (MOV) (Yang et al., 2021) 4,000 3 LtR
8.1. Qualitative Analysis
We already presented a detailed case study of CS Rankings presented as an example across the previous sections. To evaluate ShaRP  across different settings, we conducted two additional experiments. First, we analyzed a set of simple synthetic datasets coming from multiple different distributions and studied how each distribution affects the ranking. Secondly, we compared the explanations resulting from two different LtR rankers for the Moving Company dataset.
8.1.1. Score-based Ranking with Synthetic Data
In these experiments presented in more detail in Appendix B, we examine simple datasets consisting of two features and various distributions. Using this simple set-up we examine how different distributions interact with ranking. We use two settings, first, we vary the dataset distributions and fix the scoring function. Secondly, we fix the dataset distributions and vary the scoring functions.
When the scoring function is fixed but the distribution varies, feature importance depends on both the distribution and the stratum. Features with the greatest variance typically have the highest importance for ranking, with items scoring highest on these features ranking at the top. The opposite occurs at the bottom, while in the middle, either feature can dominate, increasing variance in feature importance for both features. For negatively correlated features, the same pattern emerges, but contributions have opposite signs. Discrete features, such as a Bernoulli variable, play a distinct role by segmenting the ranking: items with a value of 1 occupy the top-50%, while those with 0 fall into the bottom-50%. Within each segment, a second feature determines rank.
When we fix the data distribution but vary the ranking function, feature importance can have very different variance per stratum. This depends on the ranking function and on the feature variance. We also show that the feature with the lower variance can become the most important feature when its weight is very large.
Finally, we show that depending on the feature distribution, it is more or less possible for items ranked low to jump to the top-k when they change their values. This contradicts the assumption made in Anahideh and Mohabbati-Kalejahi (2022) where they claim that examining the local neighborhood of an item is sufficient to provide accurate feature importance explanations. We show that even items from the top-50% can move to the top-10% for example under specific distributional assumptions.
(a) XGB (original data)
(b) XGB (fairness interv.)
(c) LGB (original data)
(d) LGB (fairness interv.)
Figure 5. Feature contribution to the rank QoI for (a) XGB over the original moving company dataset, (b) XGB over the unbiased version, (c) LGB over the original moving company dataset, and (d) LGB over the unbiased version.
8.1.2. Learning to Rank
We now showcase how ShaRP  can be used to audit black-box rankers and understand their decision process. We use an XGB ranker with a pairwise ranking objective and an LGB ranker with a LambdaRank objective. Both are trained on training sets and evaluated on test sets of 2,000 tuples each. We use  ShaRP  to explain 100 items (10 per stratum) of each test set, with no approximations and the rank QoI. In Figure 5, we observe that the two LtR models behave significantly differently.
XGB rankers do not appear to rely on the Sex feature, regardless of whether the de-biasing intervention in (Yang et al., 2021) was applied. However, Race plays a significant role; in Figure 5a, Race increases the ranking of all applicants up to the 70th percentile by approximately 400 positions. This is particularly relevant as the Weight Lifting feature contributes positively to ranks in the 70-80th percentile range but negatively in the 60-70th percentile range. Ideally, its importance should be more monotonic, as seen (to some extent) in Figure 5b. While Race has slightly lower relevance overall, it remains an important feature.
In contrast, LGB rankers utilize all features in most cases. Feature importance in an LGB model trained on the original data (Figure 5c) shows that Sex and Race are highly influential across all strata, often ranking as the most important features for applicants in the lowest percentiles (50th onward). Analysis of the 10-20th, 60-70th, and 90-100th percentiles reveals that Weight Lifting has little to no impact on the model’s decisions. Thus, Sex and Race largely determine whether an applicant achieves a high rank. The fairness intervention mitigates this issue to some extent (Figure 5d) by shifting Weight Lifting’s importance toward a more intuitive and desirable role. However, Race and Sex remain highly significant, sometimes even outweighing Weight Lifting.
In summary, results indicate that XGB relies more on Race, while LGB emphasizes Sex. Bias mitigation is effective up to the 10th percentile but fails to correct bias across the remaining strata.
8.2. Comparison to Other Methods
In this section, we compare explainability methods using the metrics from Section 6, focusing on a subset of methods from Section 2. Since ShaRP and Shapley values emphasize individual explanations, we exclude methods designed for global explanations, such as those by Yang et al. and Gale and Marian (Yang et al., 2018; Gale and Marian, 2020). To compare with HIL (Yuan and Dasgupta, 2023), we adapt their code to support any score-based ranker and real data (see Appendix C), and we focus only on their weight-based methods, as their linear-approximation of Shapley values is already covered by SHAP. We do not compare to PrefShap (Hu et al., 2022), which is limited to pairwise datasets using a generalized preferential kernel function.
We compare to HRE (Anahideh and Mohabbati-Kalejahi, 2022) but use only four of their internal methods as provided by their public code base (Decision Trees (DT), Linear Regression (LR), Ordinary Least Squares (OLS), and Partial Least Squares (PLS)) and their default neighborhood settings (5-10 consecutive positions above and below the item being explained). We compare to DEXER which fits a linear regression model to the ranks and explains this model using the score-based SHAP instead of the original blackbox, treating rank as a score. Finally, we compare to SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016b), due to their wide use and availability, even though they are not designed for ranking.
8.2.1. Sensitivity
(a) HRE LR
(b) DEXER
(c) LIME
(d) SHAP
(e) ShaRP Rank
(f) HIL Std Rank
Figure 6. Sensitivity results for CS Rankings. Each dot represents a neighbor of the reference item. The x-axis shows the Euclidean explanation distance, the y-axis the absolute rank difference, and the hue indicates feature similarity to the reference item. Methods using rank as the profit function (ShaRP and HIL Std rank) perform better, with ShaRP leading. These are the only methods that consistently provide similar explanations for items that are both feature-similar and have similar outcomes.
Figure 6 presents the results comparing sensitivity for all methods. In this figure, we evaluate the similarity of explanations for pairs of similar items. For each pair of items, we compute three distances: (1) Euclidean distance between the explanations (x-axis); (2) distance between the ranks of the two items (y-axis); and (3) Euclidean distance between the items in terms of their feature values (hue, where lighter means closer). To make the plot, we place one item (the reference item) at position (0,0) and use a scatter point for each other item (neighbor), indicating the distance in ranks and the distance of the explanations. The color of the scatter point indicates the distance between the features of the reference point and the neighbor. We then overlay the plots for all items in the dataset, so that all items are used as reference points and all items are used as neighbors.
Intuitively, items with similar features and close rankings should have similar explanations. That is, the points should align near the diagonal $y=x$, with a darkening hue as the distance from the origin increases. However, this rarely holds in practice. For instance, if one feature dominates, proximity in feature space may not imply explanation similarity. Likewise, items that appear very different in feature space may still receive similar outcomes. Ideally, explanations should reflect these nuances: differing for closely ranked items with distinct features and never being too similar for items far apart in ranking. This suggests that the plot should fill the space below $y=x$, with hue darkening as the distance from the origin increases.
In Figure 6, only methods using the rank QoI produce explanations that follow the expected shape. Both ShaRP  with rank QoI (Figure 6e) and HIL-Std-Rank (Figure 6f, using our expanded definition) successfully generate similar explanations for similarly ranked items with similar features. ShaRP  performs slightly better in clustering points into denser similarity groups. In contrast, SHAP (Figure 6d), a score-based method, relies primarily on feature distances. Its plot shows distinct bands that darken away from 0 and below the diagonal. However, items ranked very differently but with similar features still receive nearly identical explanations.
LIME (Figure 6c), another score-based method, appears to incorporate both feature and ranking distance but fails to distinguish explanations as clearly as rank-based methods.
We see that score-based methods are unable to account for nonlinear relationships between score and rank. DEXER (Figure 6b), which uses linear regression to predict rank and SHAP to explain it, performs almost identically to score-based methods. While a non-linear prediction model might predict ranks more accurately, we follow a more general approach and directly incorporate ranks into the Shapley value utility function.
Finally, the explanation distance of the HRE method (Figure 6a) does not appear to depend on the ranking distance or the feature distance as we see items with similar and dissimilar explanations across all ranks and hues. This is expected, as the explanations depend only on neighbors in terms of ranking (that can have different features) and there is no guarantee that a different cluster of explanations for a further neighborhood will be different from the neighborhood of the reference point.
In Appendix D, we present sensitivity results for all methods, extensively compare ShaRP  with rank QoI to HIL with Std rank, and conduct a similar analysis for a score-based task. We show that ShaRP , which leverages Shapley values, outperforms HIL Std rank, which directly uses model weights, across all datasets. Additionally, we compare ShaRP with score QoI to ShaRP with rank QoI, demonstrating that for score-based tasks, the score QoI aligns with the diagonal line—highlighting the importance of selecting a QoI relevant to the explainability question.
8.2.2. Fidelity
It is possible to calculate Fidelity for SHAP, LIME, ShaRP, and the HIL-score. It is impossible to compute Fidelity for HIL-rank and all the HRE methods. All methods except HIL-score perform very well.
We compute the Fidelity averaged across all items in all datasets. All methods are executed using their recommended settings to compute explanations for score QoI. Additionally, we compute fidelity for ShaRP  for the rank QoI. Recall that ShaRP  is the only method that can compute a Shapley value explanation for this QoI. LIME, SHAP and ShaRP  are all achieving high explanation fidelity, on average ranging from 0.94-0.98, 0.97-1.00, and 1.00 correspondingly. HIL has reasonable fidelity for CSR (0.85) but does not perform consistently on other datasets ranging from 0.14-0.64. We include a table to visualize these results in Appendix D.
8.2.3. Agreement Between Explanations
Figure 7 presents agreement between ShaRP  and all other methods averaged across all datasets. We also provide the agreement per dataset in Appendix D. We use rank and score QoIs for this comparison, as they match those used by the methods we evaluate. Kendall’s tau distance is computed to enable cross-method comparisons. We observe that explanations vary significantly by method, regardless of the QoI. ShaRP aligns most closely with LIME and SHAP across both rank and score QoIs. HRE, which relies on localized information, naturally differs. However, even among HRE variants, explanations remain inconsistent. The two HIL methods and the two ShaRP methods produce similar explanations despite using different QoIs, suggesting that explanation agreement depends on both the QoI and the underlying explanation method In contrast, DEXER, which fits a linear regression to the ranking output and applies SHAP, differs greatly from all methods, indicating that rank cannot be effectively explained without a rank QoI.
Figure 7. Method agreement averaged across all datasets
8.3. Efficiency and Approximation
Table 2. Running time of exact computation.
avg. time (sec)
dataset # tuples # features score rank pair
ATP 86 6 0.004 0.026 0.004
CSR 189 5 0.002 0.022 0.003
THE 1,397 5 0.011 0.423 0.007
SYN 2,000 3 0.002 0.126 0.003
(a) Speedup vs. sample size, THE
(b) Speedup vs. max coalition size, THE
(c) Fidelity vs. max coalition size, THE
8.3.1. Running time of exact computation
In our first experiment, we measure the running time of the exact computation for the rank and score QoIs, and the pairwise method using the rank QoI for 3 of the real and 1 of the synthetic datasets in Table 1. We include only 1 synthetic dataset because all synthetic datasets have the same number of points ($m=2,000$) and at most 3 features, and differ in the structure of the correlations within the data. However, these correlations do not impact the running time of exact computation. Finally, we do not include the running time for the top-$k$ QoI because it is based closely on rank QoI in our implementation, and the running times of generating explanations for these two QoIs are indistinguishable.
We observe that the running time of rank and score QoIs increases with both the number of items ($m$ in Algorithm 1) and the number of features $d$. This is expected, as exact feature importance for score and rank QoIs scales linearly with dataset size $\mathcal{D}$ ($m=|\mathcal{D}-1|$ samples for exact computation) and exponentially with $d$ (due to $12^{d}-$ 1 feature coalitions). Pairwise methods, like all baseline methods, use only two data points (the pair), making their running time independent of $m$ but exponential in $d$. However, we use the pairwise method with the rank QoI which requires computing the new rank of an item or pair upon intervention in Shapley value computation (line 7 in Algorithm 2). This rank computation can be implemented in various ways. In our implementation, it scales linearly with $m$, explaining why pairwise QoI for THE ($m=1,397$, $d=5$) has a higher running time than ATP ($m=86$, $d=6$).
8.3.2. Running time and quality of approximation
We implement two methods for approximating feature importance to reduce the running time: reduce the number of samples and restrict maximum coalition size. In addition to measuring the running time, we also measure and report explanation fidelity (Eq. 8) to quantify the quality of the approximation. Figure 8 presents these results for the THE dataset, where exact computation was slowest and so there is most substantial room for improvement. In Appendix D.4 we show similar results for ATP and CSR.
Figure 8a shows the speed-up achieved by reducing the number of samples $m$. Lowering $m$ from 1378 (exact) to 20, while maintaining a maximum of size 4 coalitions (the largest possible for 5 features), accelerates rank QoI by a factor of 32—reducing runtime from 0.423 sec to 0.013 sec. Crucially, this performance gain does not compromise fidelity, which remains above 0.98 (out of 1) across all sample sizes in all experiments.
Figures 8b and 8c show the speed-up and fidelity, respectively, when coalition size is bounded. We observe that speed-up is most substantial for all QoIs for coalitions of size 1, but that fidelity is lower for this approximation method. We observe fidelity values of at least 0.83 for rank, 0.91 for score, and 0.98 for pairwise. Fidelity further improves for rank and score for coalitions of size at most 2, reaching 0.87, 0.94, and 1, respectively. The difference between the approximation methods is expected because they affect different aspects of the sampling process. Sample size affects the number of samples we draw while maximum coalition size affects the number of features we can sample concurrently.
In summary, reducing the number of samples and bounding coalition size are both effective at improving the running time, while retaining high explanation fidelity and method agreement in practice. Nonetheless, unless otherwise stated, we used exact computation in this section to reduce variability in the reported results.
9. Conclusions and Future Work
We introduced ShaRP , a comprehensive framework for quantifying feature importance in selection and ranking. Given the impact of rankers on individuals, organizations, and populations, understanding their decisions is crucial for auditing and compliance (ensuring legal adherence), recourse (helping individuals improve outcomes), and design (optimizing ranking procedures). Our work addresses the interpretability needs of these tasks.
We demonstrated the effectiveness of ShaRP  through a qualitative analysis of an impactful real-world task—the ranking of Computer Science departments. This was complemented by a thorough evaluation on real and synthetic datasets, revealing that our defined quantities of interest provide valuable and complementary insights beyond simple score-rank relationships. We showed that feature importance varies with data distribution even when the scoring function is fixed and exhibits locality. Finally, we compared ShaRP  to other local feature-based explanation methods, showing it performs favorably.
A limitation of our work is that our evaluation of the usefulness of explainability did not include a user study. Understanding how humans make sense of feature importance in ranking and act on their understanding, is in our immediate plans. We also plan to support additional quantities of interest, and develop a comprehensive benchmark that compares both local and global feature importance methods for ranking in terms of their usability, expressiveness, and performance.
Appendix A Datasets
A.1. Datasets